When an alpha particle, which has charge +2e, moves from a point with an electric potential equal to 250 V to a point with an electric potential of 1,250 V, what is the change in the electric potential energy?
Answer #1

LV= (2)(1.6×10-19)(1250-250)= 3.2×10-16J